OpenStudy (anonymous):

the graph of y = |x – 1| – 1

OpenStudy (jdoe0001):

y = |x – 1| – 1 really means y = +1(x – 1) – 1 y = -1(x – 1) – 1 graph those two, and that's your graph for the absolute value function